# Toward Stable Semi-Supervised Remote Sensing Segmentation via Co-Guidance and Co-Fusion

### v1

@@ I Introduction
Remote sensing imagery has become indispensable for a wide range of Earth observation applications, including land-cover mapping, urban planning, ecological monitoring, and disaster assessment [li2024review, huang2023deep, aleissaee2023transformers]. With the rapid advancement and widespread deployment of high-resolution satellite and aerial sensors, remote sensing data now exhibit large spatial coverage [xia2023openearthmap], complex scene layouts [wang2022unetformer], and fine-grained category structures [gao2021stransfuse]. Consequently, pixel-wise semantic segmentation has emerged as a core task, requiring models to be simultaneously sensitive to high-level semantics (e.g., category identity and contextual relationships) and low-level details (e.g., object boundaries and small structures [ma2021factseg]). Nevertheless, the fully supervised deep models that currently dominate this field heavily depend on exhaustive pixel-level annotations, which are notoriously expensive and time-consuming to obtain due to the need for expert interpretation [hua2021semantic]. In practice, only a small fraction of available images can be labeled, while the vast majority remain unlabeled [yu2023deep]. Under such conditions, there is a pressing need for semi-supervised segmentation methods that can effectively exploit abundant unlabeled data, cope with label scarcity, and maintain stable optimization despite the presence of noisy pseudo-labels [ran2024semi, zhu2025regionmatch].
To mitigate the reliance on massive annotations, recent research has made substantial strides in semi-supervised learning, evolving through three primary paradigms in the remote sensing domain: generative adversarial networks (GANs) [hung2018adversarial, souly2017semi, mittal2019semi], consistency-based methods [tarvainen2017mean, ke2019dual, chen2021semi, sohn2020fixmatch, yang2023revisiting], and pseudo-labeling methods [yang2022st++, wang2022semi, jin2024dynamic, feng2022dmt]. Early GAN-based approaches, exemplified by Hung et al. [hung2018adversarial] and Souly et al. [souly2017semi], align feature distributions via adversarial learning but often suffer from training instability and convergence difficulties. Subsequently, consistency-based methods like FixMatch [sohn2020fixmatch] and UniMatch [yang2023revisiting] became mainstream by enforcing prediction invariance under perturbations; however, they remain susceptible to confirmation bias, where incorrect predictions are reinforced over time. Similarly, pseudo-labeling strategies like U2PL [wang2022semi] and DWL [huang2024decouple] expand the training set with high-confidence predictions, yet they struggle with noise accumulation when the initial predictions are unreliable. Crucially, since these paradigms primarily rely on self-generated supervisory signals to guide training, they remain vulnerable to pseudo-label drift, particularly in complex remote sensing scenes. Without strong external guidance to rectify errors, these methods frequently encounter difficulties in distinguishing semantically similar categories or delineating precise object boundaries under severe label scarcity.
@@
    We develop a global-local feature collaborative fusion strategy, effectively fusing the global contextual information captured by CLIP with the local details produced by DINOv3. This complementary fusion enables the model to generate highly precise segmentation results.
Extensive experiments on six diverse remote sensing benchmarks, including WHDLD [shao2020multilabel], LoveDA [wang2021loveda], Potsdam [isprs-annals-I-3-293-2012], GID-15 [tong2020land], MER, and MSL [li2022stepwise], demonstrate the effectiveness of Co2S. The results show that our framework consistently delivers leading performance across various partition protocols, outperforming existing semi-supervised methods, particularly in regimes with extremely scarce annotations.
@@ II-A Remote Sensing Semi-Supervised Semantic Segmentation
In the early exploration phase, generative adversarial networks were incorporated into the Semi-Supervised segmentation framework to serve as an auxiliary supervision branch, facilitating model optimization through the adversarial discrimination of prediction distributions. Souly et al. [souly2017semi] leveraged a generator network to produce fake images using class-level information as extra training examples, forcing real samples to cluster in the feature space to improve pixel classification. Focusing on the output space, Hung et al. [hung2018adversarial] designed a discriminator in a fully convolutional manner to differentiate predicted probability maps from the ground truth, enabling the discovery of trustworthy regions in unlabeled data to facilitate effective semi-supervised learning.
Subsequently, consistency-based methods gained prominence by enforcing prediction invariance under perturbations. Meanwhile, FixMatch [sohn2020fixmatch] utilizes high-confidence predictions from weakly augmented images to supervise strongly augmented ones, providing a simplified yet effective means to leverage unlabeled data. Building on this, UniMatch [yang2023revisiting] introduces dual-stream input and feature perturbations to guide two strong views with a common weak view, successfully exploring a broader perturbation space for superior segmentation accuracy. Tailored for remote sensing, WSCL [lu2023weak] employs strong data augmentation and adaptive reweighting to decouple self-biased predictions and alleviate overfitting. Similarly, MUCA [wang2025semi] integrates multiscale uncertainty consistency and cross-teacher-student attention to construct discriminative feature representations, improving the distinction of highly similar objects. Parallel to consistency strategies, pseudo-labeling methods represent another active research direction, focusing on generating high-quality labels for self-training. U2PL [wang2022semi] enhances data utilization by identifying unreliable pixels based on entropy and using them as negative samples. Addressing the specific challenges of noise and class imbalance in remote sensing, DWL [huang2024decouple] employs a decoupled learning module alongside a ranking-based weighting module to adaptively filter and balance pseudo-labels. Different from these methods that rely on homogeneous co-training architectures or standard teacher-student paradigms, Co2S introduces a heterogeneous dual-student paradigm. By exploiting the complementary nature of distinct vision foundation models, our framework breaks the confirmation bias inherent in conventional co-training, offering an effective solution for pseudo-label drift.
@@ II-B Semantic Segmentation with Vision Foundation Models
Vision foundation models have marked a paradigm shift in semantic segmentation, branching into vision-language models (VLMs) and self-supervised models (SSMs). VLMs [radford2021learning, rao2022denseclip, zhou2022extract, liu2024remoteclip, guo2024skysense] leverage text-image alignment for transferable visual understanding. CLIP [radford2021learning] utilizes large-scale contrastive pre-training to distill generic visual concepts, while methods like DenseCLIP [rao2022denseclip] and MaskCLIP [zhou2022extract] adapt it for dense prediction tasks. In remote sensing, RemoteCLIP [liu2024remoteclip] aligns visual features of remote sensing images with textual descriptions to learn robust visual features with rich semantics. Parallelly, SSMs [caron2021emerging, oquab2023dinov2, simeoni2025dinov3, sun2022ringmo, he2022masked, peng2022beit, zhou2021ibot, xie2022simmim] focus on extracting intrinsic spatial features, capturing hierarchical visual patterns without relying on manual annotations or textual supervision. The DINO family [caron2021emerging, oquab2023dinov2, simeoni2025dinov3] employs self-distillation to learn high-quality dense features effectively, whereas RingMo [sun2022ringmo] leverages generative self-supervised learning to specifically target dense and small objects in remote sensing scenes, Cloud-Adapter [cloud-adapter] proposes a parameter-efficient tuning strategy that adapts DINOv2 [oquab2023dinov2] to the cloud segmentation without retraining the backbone.
Building on these advancements, recent studies explore architectures incorporating both VLM and SSM components [li2025exploring, zhang2025frozen, barsellotti2025talking]. Barsellotti et al. [barsellotti2025talking] proposed Talk2DINO, which aligns CLIP’s textual embeddings with DINO’s visual features via a learned mapping, resulting in less noisy segmentations with improved foreground distinction. For remote sensing, Ye et al. [ye2025towards] developed GSNet, employing query-guided feature fusion to integrate representations from generalist and specialist encoders to achieve more accurate mask predictions. Similarly, RSKT-Seg [li2025exploring] aggregates multi-directional cost maps to capture rotation-invariant cues, thereby enhancing both inference efficiency and segmentation accuracy. Alternatively, shifting the focus to utilizing frozen representations, Zhang et al. [zhang2025frozen] leveraged the static backbones of CLIP and DINO to generate high-quality pseudo-labels, enabling effective weakly supervised segmentation with reduced training costs. In a similar vein, CLIP-DINOiser [wysoczanska2024clip] applies DINO’s spatial affinities to refine the dense MaskCLIP features, producing smoother segmentation outputs without the need for pixel-level annotations. However, these approaches primarily target open-vocabulary supervised or weakly supervised scenarios. The potential of synergizing CLIP and DINO within a semi-supervised mutual learning framework, specifically to mitigate pseudo-label drift under label scarcity, remains unexplored. Our work bridges this gap by integrating their distinct priors to stabilize training when pixel-level annotations are severely limited.
@@ III-A Stable Dual-Student Semi-supervised Learning Framework
Co2S constructs a dual-student architecture comprising two heterogeneous ViT-based vision foundation models initialized from CLIP [radford2021learning] and DINOv3 [simeoni2025dinov3], respectively. As illustrated in Fig. 2, the training set consists of a labeled subset $\mathcal{D}_{l}=\{(\mathbf{x}_{l},\mathbf{y})\}$ and an unlabeled subset $\mathcal{D}_{u}=\{\mathbf{x}_{u}\}$, where images $\mathbf{x}_{l},\mathbf{x}_{u}\in\mathbb{R}^{3\times H\times W}$ and ground truth labels $\mathbf{y}\in\mathbb{R}^{1\times H\times W}$.
@@
For a labeled sample $(\mathbf{x}_{l},\mathbf{y})$, we generate a weakly augmented view $\mathbf{x}_{l}^{w}\in\mathbb{R}^{3\times H\times W}$, which is fed into both students. The CLIP-based student produces the global prediction $\mathbf{M}_{\text{pred}}^{\text{global}}\in\mathbb{R}^{1\times H\times W}$, and the DINOv3-based student produces the local prediction $\mathbf{M}_{\text{pred}}^{\text{local}}\in\mathbb{R}^{1\times H\times W}$. The supervised training objective is applied independently to each student:
$\mathcal{L}_{\text{sup}}^{\text{global}}=\mathrm{CE}(\mathbf{M}_{\text{pred}}^{\text{global}},\mathbf{y}),\quad\mathcal{L}_{\text{sup}}^{\text{local}}=\mathrm{CE}(\mathbf{M}_{\text{pred}}^{\text{local}},\mathbf{y}).$ (1)
@@
where $\text{CE}(\cdot,\cdot)$ represents the cross-entropy calculation, $\mathcal{L}_{\text{sup}}^{\text{global}}$ and $\mathcal{L}_{\text{sup}}^{\text{local}}$ denote the supervised losses for the CLIP and DINOv3 students, respectively.
@@
For an unlabeled sample $\mathbf{x}_{u}$, we strictly adhere to the dual-stream perturbation paradigm of UniMatch [yang2023revisiting]. We generate a weakly augmented view $\mathbf{x}_{u}^{w}$, two strongly augmented views $\mathbf{x}_{u}^{s1},\mathbf{x}_{u}^{s2}$, and one feature-perturbed view $\mathbf{x}_{u}^{fp}$, where $\mathbf{x}_{u}^{w},\mathbf{x}_{u}^{s1},\mathbf{x}_{u}^{s2},\mathbf{x}_{u}^{fp}\in\mathbb{R}^{3\times H\times W}$. Let $\mathbf{p}^{w}\in\mathbb{R}^{K\times H\times W}$ denote the probability map from the weak view. For each pixel at location $(i,j)$, we derive the pseudo-label $\hat{y}_{u,i,j}=\arg\max_{k}(\mathbf{p}^{w}_{i,j})$ and the confidence score $c_{i,j}=\max_{k}(\mathbf{p}^{w}_{i,j})$, where $\hat{y}_{u,i,j}\in\{1,\dots,K\}$ and $c_{i,j}\in[0,1]$. To formulate the consistency regularization, we first define a generic masked cross-entropy function $\mathcal{H}$ between a target pseudo-label map $\hat{\mathbf{y}}\in\mathbb{R}^{1\times H\times W}$ and a prediction map $\mathbf{p}\in\mathbb{R}^{K\times H\times W}$:
$\mathcal{H}(\mathbf{p},\hat{y},c)=\sum_{i,j}\mathbf{1}(c_{i,j}\geq\tau)\cdot\mathrm{CE}(\mathbf{p}_{i,j},\hat{y}_{i,j}),$ (2)
@@
where $\mathbf{1}(\cdot)$ is the indicator function and $\tau$ is the confidence threshold. The final consistency loss $\mathcal{L}_{\text{ct}}$ is calculated as a weighted sum of the losses from the feature-perturbed view and the two strongly augmented views:
$\mathcal{L}_{\text{ct}}=\frac{1}{2}\mathcal{H}(\mathbf{p}^{fp},\hat{\mathbf{y}}_{u},c)+\frac{1}{4}\big(\mathcal{H}(\mathbf{p}^{s1},\hat{\mathbf{y}}_{u},c)+\mathcal{H}(\mathbf{p}^{s2},\hat{\mathbf{y}}_{u},c)\big),$ (3)
@@
where ${\mathbf{p}^{fp},\mathbf{p}^{s1},\mathbf{p}^{s2}}\in\mathbb{R}^{K\times H\times W}$ represent the predictions resulting from the feature perturbation and the two strong augmentation views, respectively.
@@
The total training objective is formulated as the weighted sum of the supervised, consistency, and stability loss:
$\mathcal{L}=\lambda_{\text{sup}}\mathcal{L}_{\text{sup}}+\lambda_{\text{ct}}\mathcal{L}_{\text{ct}}+\lambda_{\text{sta}}\mathcal{L}_{\text{sta}},$ (4)
where we set $\lambda_{\text{sup}}=\lambda_{\text{ct}}=0.5$ in default. To prevent noisy interaction during the early training phase, $\lambda_{\text{sta}}$ follows a cosine ramp-up schedule. The detailed definition of stability loss $\mathcal{L}_{\text{sta}}$ is provided in Sec. III-C.

### v2

@@ I Introduction
Remote sensing imagery has become indispensable for a wide range of Earth observation applications, including land-cover mapping, urban planning, ecological monitoring, and disaster assessment [17, 11, 1]. With the rapid advancement and widespread deployment of high-resolution satellite and aerial sensors, remote sensing data now exhibit large spatial coverage [45], complex scene layouts [40], and fine-grained category structures [7]. Consequently, pixel-wise semantic segmentation has emerged as a core task, requiring models to be simultaneously sensitive to high-level semantics (e.g., category identity and contextual relationships) and low-level details (e.g., object boundaries and small structures [22]). Nevertheless, the fully supervised deep models that currently dominate this field heavily depend on exhaustive pixel-level annotations, which are notoriously expensive and time-consuming to obtain due to the need for expert interpretation [10]. In practice, only a small fraction of available images can be labeled, while the vast majority remain unlabeled [50]. Under such conditions, there is a pressing need for semi-supervised segmentation methods that can effectively exploit abundant unlabeled data, cope with label scarcity, and maintain stable optimization despite the presence of noisy pseudo-labels [28, 55].
To mitigate the reliance on massive annotations, recent research has made substantial strides in semi-supervised learning, evolving through three primary paradigms in the remote sensing domain: generative adversarial networks (GANs) [13, 34, 23], consistency-based methods [37, 15, 4, 33, 47], and pseudo-labeling methods [48, 42, 14, 6]. Early GAN-based approaches, exemplified by Hung et al. [13] and Souly et al. [34], align feature distributions via adversarial learning but often suffer from training instability and convergence difficulties. Subsequently, consistency-based methods like FixMatch [33] and UniMatch [47] became mainstream by enforcing prediction invariance under perturbations. However, they remain susceptible to confirmation bias, where incorrect predictions are reinforced over time. Similarly, pseudo-labeling strategies like U2PL [42] and DWL [12] expand the training set with high-confidence predictions, yet they struggle with noise accumulation when the initial predictions are unreliable. Crucially, since these paradigms primarily rely on self-generated supervisory signals to guide training, they remain vulnerable to pseudo-label drift, particularly in complex remote sensing scenes. Without strong external guidance to rectify errors, these methods frequently encounter difficulties in distinguishing semantically similar categories or delineating precise object boundaries under severe label scarcity.
@@
    We develop a global-local feature co-fusion strategy, effectively fusing the global contextual information captured by CLIP with the local texture details produced by DINOv3. This complementary fusion enables the model to generate highly precise segmentation results.
Extensive experiments on six diverse remote sensing benchmarks, including WHDLD [31], LoveDA [39], Potsdam [30], GID-15 [38], MER [18], and MSL [18], demonstrate the effectiveness of Co2S. The results show that our framework consistently delivers leading performance across various partition protocols, outperforming existing semi-supervised methods, particularly in regimes with extremely scarce annotations.
@@ II-A Remote Sensing Semi-Supervised Semantic Segmentation
In the early exploration phase, generative adversarial networks were incorporated into the Semi-Supervised segmentation framework to serve as an auxiliary supervision branch, facilitating model optimization through the adversarial discrimination of prediction distributions. Souly et al. [34] leveraged a generator network to produce fake images using class-level information as extra training examples, forcing real samples to cluster in the feature space to improve pixel classification. Focusing on the output space, Hung et al. [13] designed a discriminator in a fully convolutional manner to differentiate predicted probability maps from the ground truth, enabling the discovery of trustworthy regions in unlabeled data to facilitate effective semi-supervised learning.
Subsequently, consistency-based methods gained prominence by enforcing prediction invariance under perturbations. Meanwhile, FixMatch [33] utilizes high-confidence predictions from weakly augmented images to supervise strongly augmented ones, providing a simplified yet effective means to leverage unlabeled data. Building on this, UniMatch [47] introduces dual-stream input and feature perturbations to guide two strong views with a common weak view, successfully exploring a broader perturbation space for superior segmentation accuracy. Tailored for remote sensing, WSCL [21] employs strong data augmentation and adaptive reweighting to decouple self-biased predictions and alleviate overfitting. Similarly, MUCA [41] integrates multiscale uncertainty consistency and cross-teacher-student attention to construct discriminative feature representations, improving the distinction of highly similar objects. Parallel to consistency strategies, pseudo-labeling methods represent another active research direction, focusing on generating high-quality labels for self-training. U2PL [42] enhances data utilization by identifying unreliable pixels based on entropy and using them as negative samples. Addressing the specific challenges of noise and class imbalance in remote sensing, DWL [12] employs a decoupled learning module alongside a ranking-based weighting module to adaptively filter and balance pseudo-labels. Different from these methods that rely on homogeneous co-training architectures or standard teacher-student paradigms, Co2S introduces a heterogeneous dual-student paradigm. By exploiting the complementary nature of distinct vision foundation models, our framework breaks the confirmation bias inherent in conventional co-training, offering an effective solution for pseudo-label drift.
@@ II-B Semantic Segmentation with Vision Foundation Models
Vision foundation models have marked a paradigm shift in semantic segmentation, branching into vision-language models (VLMs) and self-supervised models (SSMs). VLMs [27, 29, 53, 19, 8] leverage text-image alignment for transferable visual understanding. CLIP [27] utilizes large-scale contrastive pre-training to distill generic visual concepts, while methods like DenseCLIP [29] and MaskCLIP [53] adapt it for dense prediction tasks. In remote sensing, RemoteCLIP [19] aligns visual features of remote sensing images with textual descriptions to learn robust visual features with rich semantics. Parallelly, SSMs [3, 24, 32, 36, 9, 25, 54, 46] focus on extracting intrinsic spatial features, capturing hierarchical visual patterns without relying on manual annotations or textual supervision. The DINO family [3, 24, 32] employs self-distillation to learn high-quality dense features effectively, whereas RingMo [36] leverages generative self-supervised learning to specifically target dense and small objects in remote sensing scenes, Cloud-Adapter [57] proposes a parameter-efficient tuning strategy that adapts DINOv2 [24] to the cloud segmentation without retraining the backbone.
Building on these advancements, recent studies explore architectures incorporating both VLM and SSM components [16, 52, 2]. Barsellotti et al. [2] proposed Talk2DINO, which aligns CLIP’s textual embeddings with DINO’s visual features via a learned mapping, resulting in less noisy segmentations with improved foreground distinction. For remote sensing, Ye et al. [49] developed GSNet, employing query-guided feature fusion to integrate representations from generalist and specialist encoders to achieve more accurate mask predictions. Similarly, RSKT-Seg [16] aggregates multi-directional cost maps to capture rotation-invariant cues, thereby enhancing both inference efficiency and segmentation accuracy. Alternatively, shifting the focus to utilizing frozen representations, Zhang et al. [52] leveraged the static backbones of CLIP and DINO to generate high-quality pseudo-labels, enabling effective weakly supervised segmentation with reduced training costs. In a similar vein, CLIP-DINOiser [44] applies DINO’s spatial affinities to refine the dense MaskCLIP features, producing smoother segmentation outputs without the need for pixel-level annotations. However, these approaches primarily target open-vocabulary supervised or weakly supervised scenarios. The potential of synergizing CLIP and DINO within a semi-supervised mutual learning framework, specifically to mitigate pseudo-label drift under label scarcity, remains unexplored. Our work bridges this gap by integrating their distinct priors to stabilize training when pixel-level annotations are severely limited.
@@ III-A Stable Dual-Student Semi-supervised Learning Framework
As illustrated in Fig. 2, Co2S adopts a dual-student semi-supervised segmentation pipeline that leverages heterogeneous priors from CLIP [27] and DINOv3 [32]. The two students process the same augmented views in parallel and output segmentation logits. Specifically, the CLIP student is guided by explicit class-level semantics encoded from text, while the DINOv3 student is driven by implicit class-level semantics represented by learnable queries. Both are injected into isomorphic yet parameter-independent query-based decoders (see Sec. III-B). For labeled data, each student is supervised by the ground-truth mask. For unlabeled data, each student is trained with weak-to-strong consistency regularization across its own augmented predictions, while a global-local collaborative fusion mechanism is further introduced to perform confidence-driven pixel-wise mutual supervision between the two students, forming a stability constraint that suppresses pseudo-label drift (see Sec. III-C).
@@
Formally, the training set consists of a labeled subset $\mathcal{D}_{l}=\{(\mathbf{x}_{l},\mathbf{y}_{l})\}$ and an unlabeled subset $\mathcal{D}_{u}=\{\mathbf{x}_{u}\}$, where $\mathbf{x}_{l},\mathbf{x}_{u}\in\mathbb{R}^{3\times H\times W}$ and $\mathbf{y}_{l}\in\mathbb{R}^{1\times H\times W}$. Each student outputs segmentation logits with $N$ classes. For a labeled sample $(\mathbf{x}_{l},\mathbf{y}_{l})$, we generate a weakly augmented view $\mathbf{x}_{l}^{w}$ and feed it into both students, obtaining $\mathbf{Z}^{\text{global}}\in\mathbb{R}^{N\times H\times W}$ from the CLIP student and $\mathbf{Z}^{\text{local}}\in\mathbb{R}^{N\times H\times W}$ from the DINOv3 student. The supervised objective is applied independently:
$\mathcal{L}_{\text{sup}}^{\text{global}}=\mathrm{CE}(\mathbf{Z}^{\text{global}},\mathbf{y}_{l}),\quad\mathcal{L}_{\text{sup}}^{\text{local}}=\mathrm{CE}(\mathbf{Z}^{\text{local}},\mathbf{y}_{l}),$ (1)
@@
where $\mathrm{CE}(\cdot,\cdot)$ denotes the pixel-wise cross-entropy between logits and the ground-truth mask, $\mathcal{L}_{\text{sup}}^{\text{global}}$ and $\mathcal{L}_{\text{sup}}^{\text{local}}$ denote the supervised losses for the CLIP and DINOv3 students, respectively.
@@
For an unlabeled sample $\mathbf{x}_{u}$, we strictly adhere to the dual-stream perturbation paradigm of UniMatch [47]. We generate a weakly augmented view $\mathbf{x}_{u}^{w}$, two strongly augmented views $\mathbf{x}_{u}^{s1},\mathbf{x}_{u}^{s2}$, and one feature-perturbed view $\mathbf{x}_{u}^{fp}$, where $\mathbf{x}_{u}^{w},\mathbf{x}_{u}^{s1},\mathbf{x}_{u}^{s2},\mathbf{x}_{u}^{fp}\in\mathbb{R}^{3\times H\times W}$. For each student, let $\mathbf{Z}^{w}\in\mathbb{R}^{N\times H\times W}$ denote the segmentation logits predicted from the weak view, and $\mathbf{p}^{w}=\mathrm{Softmax}(\mathbf{Z}^{w})$ denote the probability map, where $\mathrm{Softmax}(\cdot)$ is applied over the class dimension. For each pixel at location $(i,j)$, we assign the pseudo-label $\hat{y}_{u,i,j}=\arg\max_{n}(\mathbf{p}^{w}_{n,i,j})$ and the confidence score $c_{i,j}=\max_{n}(\mathbf{p}^{w}_{n,i,j})$, where $\hat{y}_{u,i,j}\in\{1,\dots,N\}$ and $c_{i,j}\in[0,1]$. To formulate the consistency regularization, we define a generic masked cross-entropy function $\mathcal{H}$ given a logit map $\mathbf{Z}\in\mathbb{R}^{N\times H\times W}$, a pseudo-label map $\hat{\mathbf{y}}_{u}\in\mathbb{R}^{1\times H\times W}$, and confidence scores $\mathbf{c}\in\mathbb{R}^{1\times H\times W}$.
$\mathcal{H}(\mathbf{Z},\hat{\mathbf{y}}_{u},\mathbf{c})=\sum_{i,j}\mathbf{1}(c_{i,j}\geq\tau)\cdot\mathrm{CE}(\mathbf{Z}_{:,i,j},\hat{y}_{u,i,j}),$ (2)
@@
where $\mathbf{1}(\cdot)$ is the indicator function and $\tau$ is the confidence threshold. The final consistency loss $\mathcal{L}_{\text{ct}}$ is calculated as:
$\mathcal{L}_{\text{ct}}=\frac{1}{2}\mathcal{H}(\mathbf{Z}^{fp},\hat{\mathbf{y}}_{u},\mathbf{c})+\frac{1}{4}\Big(\mathcal{H}(\mathbf{Z}^{s1},\hat{\mathbf{y}}_{u},\mathbf{c})+\mathcal{H}(\mathbf{Z}^{s2},\hat{\mathbf{y}}_{u},\mathbf{c})\Big),$ (3)
@@
where $\mathbf{Z}^{fp},\mathbf{Z}^{s1},\mathbf{Z}^{s2}\in\mathbb{R}^{N\times H\times W}$ are the logits predicted from the feature-perturbed view and the two strongly augmented views, respectively.
@@
For each student, the total training objective is formulated as the weighted sum of the supervised, consistency, and stability loss:
$\mathcal{L}=\lambda_{\text{sup}}\mathcal{L}_{\text{sup}}+\lambda_{\text{ct}}\mathcal{L}_{\text{ct}}+\lambda_{\text{sta}}\mathcal{L}_{\text{sta}}.$ (4)
Here, $\mathcal{L}_{\text{sup}}$ refers to $\mathcal{L}_{\text{sup}}^{\text{global}}$ for the CLIP student and $\mathcal{L}_{\text{sup}}^{\text{local}}$ for the DINOv3 student, $\mathcal{L}_{\text{ct}}$ and $\mathcal{L}_{\text{sta}}$ are defined analogously. We set $\lambda_{\text{sup}}=\lambda_{\text{ct}}=0.5$, and use a cosine ramp-up schedule for $\lambda_{\text{sta}}$. The detailed definition of stability loss $\mathcal{L}_{\text{sta}}$ is provided in Sec. III-C.